# PathMoCo: A Novel Framework to Improve Feature Embedding in Self-supervised Contrastive Learning for Histopathological Images

### v1

@@ Abstract
Self-supervised contrastive learning has become a cornerstone in various areas, particularly histopathological image analysis. Image augmentation plays a crucial role in self-supervised contrastive learning, as it generates variations in image samples. However, traditional image augmentation techniques often overlook the unique characteristics of histopathological images. In this paper, we propose a new histopathology-specific image augmentation method called stain reconstruction augmentation (SRA). We integrate our SRA with MoCo v3, a leading model in self-supervised contrastive learning, along with our additional contrastive loss terms, and call the new model PathMoCo. We demonstrate that our PathMoCo always outperforms the standard MoCo v3 across various downstream tasks and achieves comparable or superior performance to other foundation models pre-trained on significantly larger histopathology datasets.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author. Email: bodong.zhang@utah.edu
@@ 1 Introduction
Hematoxylin & Eosin (H&E) stained histopathological image is the dominant image type in histopathology. Hematoxylin stains the cell nuclei a deep blue or purple, while Eosin stains the cytoplasm and extracellular matrix pink, allowing for clear differentiation between different cellular and tissue components. The variations of stains across images are shown in fig. 1
@@
In this paper, we propose a novel histopathology-specific image augmentation method called stain reconstruction augmentation (SRA) and apply it to H&E stained histopathological images. We first performed stain separation [6] to decompose RGB images into Hematoxylin stain channel images and Eosin stain channel images in Optical Density (OD) space. Instead of multiplying a preset random scalar to each channel for augmentation [7, 8, 9], where the augmented images are statistically affected by the intensities of original stain channel images, we first normalize each single stain channel image by dividing it by its max intensity, then multiplying the random scalar we predefined to augment each stain channel image. By this way, the intensities of augmented stain channel images are not affected by the max intensity of original stain channel images. Additionally, when we reconstruct augmented images from OD space back to RGB space, we define a probability for excluding one stain channel and only allow the other channel to remain to provide stronger augmentations. We integrate our SRA into MoCo v3 [10], one of the state-of-the-art self-supervised contrastive learning models, as well as our additional contrastive loss terms, to form our new model-PathMoCo.
@@
  • We conducted experiments on multiple publicly available datasets. We demonstrate that our proposed PathMoCo always outperforms MoCo v3 on various downstream tasks. We also achieved comparable or superior performance to other foundation models pre-trained on significantly larger histopathology datasets.
  • We made the PathMoCo code publicly available at github.com/hamidmanoochehri/Paper_PathMoCo
Figure 2: The pipeline of our PathMoCo. We integrate our Stain Reconstruction Augmentation (SRA) as well as additional contrastive loss terms (CL3 and CL4) into MoCo v3.
@@ 3 Methods
3.1 An overview of PathMoCo
fig. 2 shows the overall workflow of PathMoCo. First, all pixels within the tissue regions of an H&E whole slide image (WSI) are collected to analyze the max intensity (strength) of each stain in the current WSI, where the intensity is measured on Optical Density (OD) space after stain separation process. To perform stain reconstruction augmentation, we predefine an absolute range for the target strength of each stain and map the real strength of each stain to a random value in this target range. Unlike the approach in [7, 8, 9], which only slightly adjusts the strength of each stain within a relative range by multiplying a random factor between 0.95 and 1.05, our PathMoCo directly defines a much broader absolute range for target strength of each stain channel. For instance, if the target range is set between 0.5 and 2, and the strength of a particular stain in a WSI is 2, this indicates a deeply stained image. Augmenting this stain channel only makes the new maximum intensity fall between 0.5 and 2, without surpassing the original strength. This method allows for more extensive and stronger augmentations while ensuring the strength remains within an appropriate range. Moreover, inspired by multi-modal contrastive learning [30, 31], we define specific probabilities for excluding one stain channel, allowing only the other stain channel to remain after augmentation. Examples of images resulting from our stain reconstruction augmentations, including single stain channel images, can be seen in fig. 3.
@@ 4.2 Experiment settings
After applying stain reconstruction augmentation, we followed the standard MoCo v3 training procedure, including additional general augmentations, with ResNet50 [38] as backbone. The encoders consistently received 224x224 patches as inputs except for PathDino, where 512x512 patches were used after image resizing. We also included an option in our code to add the augmentation contrastive loss, $CL_{aug}$. Moreover, we tracked the contrastive loss on the validation set and selected the training epoch with the lowest contrastive loss on the validation set. In the downstream tasks on the TCGA KIRC and Utah KIRC, we used a balanced sampler for training and balanced accuracy for evaluation, as the patch labels were highly imbalanced. For the downstream tasks on the Camelyon 16 dataset, we employed DTFD-MIL [37], as it has been shown to be one of the best multi-instance learning models for Camelyon 16 and other datasets. Stain reconstruction augmentation was only applied during PathMoCo pre-training, not in any downstream tasks, where the MoCo v3’s image classification code was used.
@@ 4.3 Pre-training and downstream task on same datasets
We first evaluated PathMoCo by performing contrastive feature representation learning and downstream classification on the same datasets (TCGA KIRC or Utah KIRC). Since there is no domain shift in this scenario, we consistently used the first option’s target range for stain strengths, which has narrower limits (H: [0.5, 2.0], E: [0.2, 2.0]), and set the probability of generating pure Hematoxylin or pure Eosin images to zero. After pre-training, we froze the pre-trained encoders and performed classification on the same dataset. For reference, we also compared the classification results with other state-of-the-art foundation encoder models, including PathDino pre-trained on 11,765 TCGA slides by [21], as well as Barlow Twins, SwAV, and MoCo v2 pre-trained on 20,994 TCGA slides and 15672 from the TULIP dataset by [39]. Additionally, we included the basic ResNet50 model in a fully supervised learning setting for comparison.
As shown in table 3 and table 4, our PathMoCo significantly outperforms standard MoCo v3 and stain augmentations from [7, 8]. On the TCGA KIRC dataset, PathMoCo surpasses MoCo v3 by 4.25% (83.62% vs. 79.37%) in balanced accuracy, and SRA also outperforms the traditional stain augmentation (TSA) [7, 8] by 2.12% (83.62% vs. 81.50%). On the Utah KIRC dataset, PathMoCo exceeds MoCo v3 by 2.08% (95.85% vs. 93.77%), and MoCo v3 plus TSA by 1.85% (95.85% vs. 94.00%). The comparison with other foundation models further demonstrates the advantages of PathMoCo. Despite these state-of-the-art models being pre-trained on much larger datasets, PathMoCo still reaches the best.
@@
PathMoCo 83.62 $\pm$ 0.28
Table 3: Performance of pre-trained models on TCGA KIRC dataset compared to foundation models (20X magnification). TSA represents traditional stain augmentation methods. [7, 8]
@@ 4.4 Pre-training and downstream task on different datasets
Although our PathMoCo demonstrates outstanding results as discussed in the previous subsection, the more common application of self-supervised learning is to pre-train an encoder on large datasets and then apply the frozen encoder to other datasets. To evaluate PathMoCo in a more practical setting, we pre-trained PathMoCo on 300 training slides from the TCGA KIRC dataset and subsequently evaluated the encoders on the Utah KIRC dataset and the Camelyon 16 dataset separately.
@@
Based on table 5, on the Utah KIRC dataset, we observed 2.8% improvement with PathMoCo compared to MoCo v3. Both PathMoCo and MoCo v3 pre-trained on the TCGA KIRC dataset outperform those pre-trained on the Utah KIRC dataset itself. The most likely reason is that the TCGA KIRC dataset contains significantly more patches. Even though domain shift is present, both datasets focus on kidney cancer, which mitigates the impact. When comparing the results with foundation models, it becomes evident that pre-training and downstream classification both on kidney images offer distinct advantages.
@@
Table 6: Ablation study results showing the impact of each component of PathMoCo on Utah KIRC dataset and Camelyon 16 dataset.
The multi-instance learning results on the Camelyon 16 dataset can also be found in table 5. The MoCo v3 encoder pre-trained on the TCGA KIRC dataset shows lower performance compared to encoders pre-trained on ImageNet. One possible explanation is the substantial domain shift between kidney cancer and breast cancer slides. This is further illustrated in fig. 5, which highlights the significant differences between TCGA kidney cancer patches and Camelyon 16 breast cancer patches. However, with our stain reconstruction augmentation and our additional augmentation contrastive loss, PathMoCo improves the balanced accuracy by 8.53% compared to MoCo v3, while MoCo v3 + TSA exceeds MoCo v3 by only 1.82%. When comparing results with other foundation models pre-trained on much larger datasets with a greater variety of organ types, PathMoCo surpasses Barlow Twins and achieves comparable F1 score, accuracy, and balanced accuracy to PathDino, SwAV, and MoCo v2. Our pre-trained PathMoCo encoder will be released upon publication.
@@ 4.5 Ablation studies
We also conducted ablation studies to carefully analyze the impact of each component in PathMoCo.
The ablation studies for transfer learning from the TCGA KIRC dataset to the Utah KIRC dataset are presented in table 6. From the results, we found that simply adding the augmentation contrastive loss to MoCo v3 does not yield any improvement. However, this loss becomes effective when combined with stain reconstruction augmentation. Using a wider target range for stain strength and incorporating the possibility of generating pure Hematoxylin or pure Eosin images in stain reconstruction augmentation also proved beneficial. During transfer learning, the slides exhibit significantly more variation across datasets, and PathMoCo addresses this by generating highly augmented images that are also clinically meaningful.
Lastly, we evaluated the contribution of each component in PathMoCo through downstream tasks on the Camelyon 16 dataset. We observed that both the augmentation contrastive loss and stain reconstruction augmentation independently improve performance. When combining the augmentation contrastive loss with stain reconstruction augmentation using a wider target range of stain variation, we achieved a 4.43% improvement in balanced accuracy. Unlike transfer learning from the TCGA KIRC to the Utah KIRC dataset, where setting the probability of converting augmented images to pure Hematoxylin or pure Eosin to 0.1 resulted in a minor boost, we observed a 4.16% increase in balanced accuracy on Camelyon 16 by simply adjusting this probability from 0 to 0.1, which implies that stronger augmentations are more beneficial on cases with stronger domain shift. All the adjustments contributed to PathMoCo outperforming MoCo v3 by 8.58%.
@@ 5 Conclusion
In this paper, we propose a novel framework for contrastive representation learning in pathology images. Specifically, we introduce a new stain reconstruction augmentation method and an augmentation contrastive loss. Experiments on downstream tasks, including image classification and multi-instance learning, demonstrate significant improvements when applying our algorithms to MoCo v3. Despite being pre-trained on much smaller datasets, PathMoCo achieves comparable or superior performance in various self-supervised learning and downstream tasks compared to other state-of-the-art foundational encoders.
@@ Acknowledgements
The work was funded in part by NIH 1R21CA277381. The Computational Oncology Research Initiative (CORI) at the Huntsman Cancer Institute, The Department of Pathology at the University of Utah and ARUP Laboratories had also supported the work. Part of the results presented in this paper are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga.

### v3

# SRA: A Novel Method to Improve Feature Embedding in Self-supervised Learning for Histopathological Images

@@ Abstract
Self-supervised learning has become a cornerstone in various areas, particularly histopathological image analysis. Image augmentation plays a crucial role in self-supervised learning, as it generates variations in image samples. However, traditional image augmentation techniques often overlook the unique characteristics of histopathological images. In this paper, we propose a new histopathology-specific image augmentation method called stain reconstruction augmentation (SRA). We integrate our SRA with MoCo v3, a leading model in self-supervised contrastive learning, along with our additional contrastive loss terms, and call the new model SRA-MoCo v3. We demonstrate that our SRA-MoCo v3 always outperforms the standard MoCo v3 across various downstream tasks and achieves comparable or superior performance to other foundation models pre-trained on significantly larger histopathology datasets.
footnotetext: *The first two authors contributed equally to this work.footnotetext: †Corresponding author. Email: bodong.zhang@utah.edu
@@ 1 Introduction
Hematoxylin & Eosin (H&E) stained histopathological image is the dominant image type in histopathology. Hematoxylin stains the cell nuclei a deep blue or purple, while Eosin stains the cytoplasm and extracellular matrix pink, allowing for clear differentiation between different cellular and tissue components. The variations of stains across images are shown in fig. 1.
@@
In this paper, we propose a novel histopathology-specific image augmentation method called stain reconstruction augmentation (SRA) and apply it to H&E stained histopathological images. We first performed stain separation [27] to decompose RGB images into Hematoxylin stain channel images and Eosin stain channel images in Optical Density (OD) space. Instead of multiplying a preset random scalar to each channel for augmentation [29, 30, 7], where the augmented images are statistically affected by the intensities of original stain channel images, we first normalize each single stain channel image by dividing it by its max intensity, then multiplying the random scalar we predefined to augment each stain channel image. By this way, the intensities of augmented stain channel images are not affected by the max intensity of original stain channel images. Additionally, when we reconstruct augmented images from OD space back to RGB space, we define a probability for excluding one stain channel and only allow the other channel to remain to provide stronger augmentations. We integrate our SRA into MoCo v3 [11], one of the state-of-the-art self-supervised contrastive learning models, as well as our additional contrastive loss terms, to form our new model SRA-MoCo v3.
@@
  • We conducted experiments on multiple publicly available datasets. We demonstrate that our proposed SRA-MoCo v3 always outperforms MoCo v3 on various downstream tasks. We also achieved comparable or superior performance to other foundation models pre-trained on significantly larger histopathology datasets.
  • We made the code publicly available at github.com/hamidmanoochehri/Paper_SRA
Figure 2: The pipeline of our SRA-MoCo v3. We integrate our Stain Reconstruction Augmentation (SRA) as well as additional contrastive loss terms (CL3 and CL4) into MoCo v3.
@@ 3 Methods
3.1 An overview of SRA-MoCo v3
fig. 2 shows the overall workflow of SRA-MoCo v3. First, all pixels within the tissue regions of an H&E whole slide image (WSI) are collected to analyze the max intensity (strength) of each stain in the current WSI, where the intensity is measured on Optical Density (OD) space after stain separation process. To perform stain reconstruction augmentation, we predefine an absolute range for the target strength of each stain and map the real strength of each stain to a random value in this target range. Unlike the approach in [29, 30, 7], which only slightly adjusts the strength of each stain within a relative range by multiplying a random factor between 0.95 and 1.05, our SRA directly defines a much broader absolute range for target strength of each stain channel. For instance, if the target range is set between 0.5 and 2, and the original strength of a particular stain in a WSI is 2, this indicates a deeply stained image. Augmenting this stain channel only makes the new maximum intensity fall between 0.5 and 2, without surpassing the original strength. While the traditional augmentation makes the new maximum intensity fall between 1.9 and 2.1. Our method allows for more extensive and stronger augmentations while ensuring the strength remains within an appropriate range. Moreover, inspired by multi-modal contrastive learning [6, 36], we define specific probabilities for excluding one stain channel, allowing only the other stain channel to remain after augmentation. Examples of images resulting from our stain reconstruction augmentations, including single stain channel images, can be seen in fig. 3.
@@ 4.2 Experiment settings
After applying stain reconstruction augmentation, we followed the standard MoCo v3 training procedure, including additional general augmentations, with ResNet50 [18] as backbone. The encoders consistently received 224x224 patches as inputs except for PathDino, where 512x512 patches were used after image resizing. We also included an option in our code to add the augmentation contrastive loss, $CL_{aug}$. Moreover, we tracked the contrastive loss on the validation set and selected the training epoch with the lowest contrastive loss on the validation set. In the downstream tasks on the TCGA KIRC and Utah KIRC, we used a balanced sampler for training and balanced accuracy for evaluation, as the patch labels were highly imbalanced. For the downstream tasks on the Camelyon 16 dataset, we employed DTFD-MIL [38], as it has been shown to be one of the best multi-instance learning models for Camelyon 16 and other datasets. Stain reconstruction augmentation was only applied during pre-training, not in any downstream tasks, where the MoCo v3’s image classification code was used.
@@ 4.3 Pre-training and downstream task on same datasets
We first evaluated SRA-MoCo v3 by performing contrastive feature representation learning and downstream classification on the same datasets (TCGA KIRC or Utah KIRC). Since there is no domain shift in this scenario, we consistently used the first option’s target range for stain strengths, which has narrower limits (H: [0.5, 2.0], E: [0.2, 2.0]), and set the probability of generating pure Hematoxylin or pure Eosin images to zero. After pre-training, we froze the pre-trained encoders and performed classification on the same dataset. For reference, we also compared the classification results with other state-of-the-art foundation encoder models, including PathDino pre-trained on 11,765 TCGA slides by [2], as well as Barlow Twins, SwAV, and MoCo v2 pre-trained on 20,994 TCGA slides and 15672 from the TULIP dataset by [21]. Additionally, we included the basic ResNet50 model in a fully supervised learning setting for comparison.
As shown in table 3 and table 4, our SRA-MoCo v3 significantly outperforms standard MoCo v3 and stain augmentations from [29, 30]. On the TCGA KIRC dataset, SRA-MoCo v3 surpasses MoCo v3 by 4.25% (83.62% vs. 79.37%) in balanced accuracy, and SRA also outperforms the traditional stain augmentation (TSA) [29, 30] by 2.12% (83.62% vs. 81.50%). On the Utah KIRC dataset, SRA-MoCo v3 exceeds MoCo v3 by 2.08% (95.85% vs. 93.77%), and MoCo v3 plus TSA by 1.85% (95.85% vs. 94.00%). The comparison with other foundation models further demonstrates the advantages of SRA-MoCo v3. Despite these state-of-the-art models being pre-trained on much larger datasets, SRA-MoCo v3 still reaches the best.
@@
SRA-MoCo v3 83.62 $\pm$ 0.28
Table 3: Performance of pre-trained models on TCGA KIRC dataset compared to foundation models (20X magnification). TSA represents traditional stain augmentation methods [29, 30]
@@ 4.4 Pre-training and downstream task on different datasets
Although our SRA-MoCo v3 demonstrates outstanding results as discussed in the previous subsection, the more common application of self-supervised learning is to pre-train an encoder on large datasets and then apply the frozen encoder to other datasets. To evaluate SRA-MoCo v3 in a more practical setting, we pre-trained SRA-MoCo v3 on 300 training slides from the TCGA KIRC dataset and subsequently evaluated the encoders on the Utah KIRC dataset and the Camelyon 16 dataset separately.
@@
Based on table 5, on the Utah KIRC dataset, we observed 2.8% improvement with SRA-MoCo v3 compared to MoCo v3. Both SRA-MoCo v3 and MoCo v3 pre-trained on the TCGA KIRC dataset outperform those pre-trained on the Utah KIRC dataset itself. The most likely reason is that the TCGA KIRC dataset contains significantly more patches. Even though domain shift is present, both datasets focus on kidney cancer, which mitigates the impact. When comparing the results with foundation models, it becomes evident that pre-training and downstream classification both on kidney images offer distinct advantages.
@@
Table 6: Ablation study results showing the impact of each component of SRA-MoCo v3 on Utah KIRC dataset and Camelyon 16 dataset.
The multi-instance learning results on the Camelyon 16 dataset can also be found in table 5. The MoCo v3 encoder pre-trained on the TCGA KIRC dataset shows lower performance compared to encoders pre-trained on ImageNet. One possible explanation is the substantial domain shift between kidney cancer and breast cancer slides. This is further illustrated in fig. 5, which highlights the significant differences between TCGA kidney cancer patches and Camelyon 16 breast cancer patches. However, with our stain reconstruction augmentation and our additional augmentation contrastive loss, SRA-MoCo v3 improves the balanced accuracy by 8.53% compared to MoCo v3, while MoCo v3 + TSA exceeds MoCo v3 by only 1.82%. When comparing results with other foundation models pre-trained on much larger datasets with a greater variety of organ types, SRA-MoCo v3 surpasses Barlow Twins and achieves comparable F1 score, accuracy, and balanced accuracy to PathDino, SwAV, and MoCo v2. Our pre-trained SRA-MoCo v3 encoder will be released upon publication.
@@ 4.5 Ablation studies
We also conducted ablation studies to carefully analyze the impact of each component in SRA-MoCo v3.
The ablation studies for transfer learning from the TCGA KIRC dataset to the Utah KIRC dataset are presented in table 6. From the results, we found that simply adding the augmentation contrastive loss to MoCo v3 does not yield any improvement. However, this loss becomes effective when combined with stain reconstruction augmentation. Using a wider target range for stain strength and incorporating the possibility of generating pure Hematoxylin or pure Eosin images in stain reconstruction augmentation also proved beneficial. During transfer learning, the slides exhibit significantly more variation across datasets, and SRA-MoCo v3 addresses this by generating highly augmented images that are also clinically meaningful.
Lastly, we evaluated the contribution of each component in SRA-MoCo v3 through downstream tasks on the Camelyon 16 dataset. We observed that both the augmentation contrastive loss and stain reconstruction augmentation independently improve performance. When combining the augmentation contrastive loss with stain reconstruction augmentation using a wider target range of stain variation, we achieved a 4.43% improvement in balanced accuracy. Unlike transfer learning from the TCGA KIRC to the Utah KIRC dataset, where setting the probability of converting augmented images to pure Hematoxylin or pure Eosin to 0.1 resulted in a minor boost, we observed a 4.16% increase in balanced accuracy on Camelyon 16 by simply adjusting this probability from 0 to 0.1, which implies that stronger augmentations are more beneficial on cases with stronger domain shift. All the adjustments contributed to SRA-MoCo v3 outperforming MoCo v3 by 8.58%.
@@ 5 Conclusion
In this paper, we propose a novel framework for contrastive representation learning in pathology images. Specifically, we introduce a new stain reconstruction augmentation method and an augmentation contrastive loss. Experiments on downstream tasks, including image classification and multi-instance learning, demonstrate significant improvements when applying our algorithms to MoCo v3. Despite being pre-trained on much smaller datasets, SRA-MoCo v3 achieves comparable or superior performance in various self-supervised learning and downstream tasks compared to other state-of-the-art foundational encoders.
@@ Acknowledgements
The work was funded in part by NIH 1R21CA277381. The Computational Oncology Research Initiative (CORI) at the Huntsman Cancer Institute, The Department of Pathology at the University of Utah and ARUP Laboratories had also supported the work. Part of the results presented in this paper are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga